# Bias in the tensor-to-scalar ratio from self-interacting dark radiation

### v1

@@ I Introduction
It is well known that the coupling between ultra-relativistic (UR) species present in the primordial plasma and the GW modifies the primordial $B$-mode spectrum of the CMB polarization [5, 6], thereby affecting the observable tensor-to-scalar-ratio today. UR species include photons and those non-photonic relativistic components commonly grouped as neutrinos and possible beyond-Standard-Model relics, referred to as dark radiation (DR). After decoupling from the plasma, UR species are typically treated as free-streaming: in that limit they develop a nonvanishing anisotropic stress which back-reacts on the GW evolution producing a damping of the amplitude of the GW spectrum. Conversely, when interactions among UR species are efficient they isotropize the distribution function, suppress anisotropic stress and thus remove the GW damping. The competition between the species’ interaction rate and the Hubble expansion rate therefore determines whether the species behave as a free-streaming gas or as a fluid-like component, with direct consequences for the amplitude and shape of the primordial GW spectrum [7, 8, 9, 10]. The current GW spectrum including these effects was explicitly computed in [8, 10], while the effect on the primordial $B$-mode spectrum is studied in this paper.
The key question motivating this work is whether self-interactions among UR species, and in particular among a cosmologically relevant DR component, can alter the observed $B$-mode signal sufficiently to bias the inference of $r$. While current CMB and large-scale-structure measurements place stringent limits on many classes of neutrino interactions at the epochs probed by the primary CMB anisotropies [11], models of new self-interacting DR remain phenomenologically viable. In these scenarios the DR abundance, parametrized by $\Delta N_{\rm eff}$, is tightly constrained by data [12, 13, 14], whereas the interaction strength can still span a wide range without being excluded [15].
@@
The current experimental landscape of CMB polarization includes imaging telescopes such as CLASS [16], ACT [17], and SPT-3G [18], which are already delivering high-quality data. QUBIC [19, 20], as the first bolometric interferometer telescope, provides a complementary approach to these measurements. Upcoming experiments, such as the Simons Observatory (SO) and the LiteBIRD satellite, are designed to reach sensitivities to primordial $B$-modes at the level of $r\sim 10^{-3}$. SO will deploy three Small Aperture Telescopes (SATs) and one Large Aperture Telescope (LAT) in the Atacama Desert, with a total of nearly 60,000 TES bolometers spanning six frequencies from 27 to 280 GHz [21]. The SATs will survey approximately 10% of the sky with a projected white noise level of $2\,\mu{\rm K}$-arcmin in combined 93/145 GHz bands, enabling a constraint on the tensor-to-scalar ratio with an expected sensitivity of $\sigma(r)\simeq 0.003$. The LAT will cover about 40% of the sky with arcminute resolution and $6\,\mu{\rm K}$-arcmin noise in the same bands. Complementing these ground-based measurements, LiteBIRD (scheduled for launch to L2 around 2032) will perform a full-sky survey over three years with 15 frequency bands between 34 and 448 GHz. Its typical angular resolution is $\sim 0.5^{\circ}$ at 100 GHz, with a target sensitivity of $2.2\,\mu{\rm K}$-arcmin in polarization, designed to achieve $\delta r\simeq 0.001$ including systematic uncertainties [22]. Together, the deep and high-resolution observations of SO and the all-sky coverage and systematic control of LiteBIRD will provide a decisive test of inflationary models. Beyond SO and LiteBIRD, the Probe of Inflation and Cosmic Origins (PICO) is a NASA probe-class mission concept designed to map the full sky with unprecedented sensitivity ($0.65\,\mu{\rm K}$-arcmin) across 21–799 GHz, enabling constraints on inflationary physics, neutrino masses, structure formation, and Galactic foregrounds [23, 24].
@@ II.1 First-order perturbations
The first-order Einstein–Boltzmann system couples the metric to the energy–momentum tensor through Einstein equations, and the one-particle distribution function (1pdf) of the different species to the collision integral through the Boltzmann equations. Since our main interest is the impact on primordial CMB $B$-modes, it is sufficient to focus only on the evolution of the tensor sector within the scalar–vector–tensor (SVT) decomposition of cosmological perturbations, assuming a homogeneous and isotropic Friedmann–Lemaître–Robertson–Walker (FLRW) background [29, 30, 3, 1].
@@
For the Boltzmann equations, we include the dynamics of the one-particle distribution function (1pdf) of DR can be written as
@@
Restricting to the tensor sector ($m=2$), the Boltzmann hierarchy becomes
$\Theta^{\prime(2)}_{2}$ $=-\frac{\sqrt{5}}{7}\,k\,\Theta^{(2)}_{3}-\alpha_{2}a\langle\Gamma\rangle\Theta^{(2)}_{2}-\dot{h},$ (14)
$\Theta^{\prime(2)}_{\ell}$ $=k\left[\frac{s_{\ell}}{2\ell+1}\Theta^{(2)}_{\ell-1}-\frac{s_{\ell+1}}{2\ell+3}\Theta^{(2)}_{\ell+1}\right]-\alpha_{\ell}a\langle\Gamma\rangle\Theta^{(2)}_{\ell},$
with $s_{\ell}=\sqrt{(\ell^{2}-4)/\ell^{2}}$. The hierarchy is closed using the standard truncation described in [33], as implemented in CLASS. The relaxation-time coefficients $\alpha_{\ell}$ were computed in [15]; for simplicity, we adopt $\alpha_{\ell}=\alpha_{2}\simeq 0.188$.
@@
Finally, we implement Eqs. (11), (14), and (15) into the Einstein–Boltzmann solver CLASS [25, 26] [^1] In practice, CLASS formulates the equations using a different variable, $F^{(m)}_{\ell}$, to describe 1pdf perturbations. The relation between $F^{(m)}_{\ell}$ and $\Theta^{(m)}_{\ell}$ is detailed in [32]. to compute the evolution of cosmological perturbations.
@@ III Impact on $C^{BB}_{\ell}$
The main effect of these interactions is an effective enhancement of the $B$-mode power spectrum amplitude across all multipoles in comparison to the standard free-streaming DR case, with the exception of $\ell\sim 100$. This enhancement grows with increasing coupling constant $\lambda$ (or, equivalently, higher $z_{\rm int}$). Earlier transitions (larger $\lambda$ or $z_{\rm int}$) correspond to stronger self-interactions that drive the 1pdf of DR toward equilibrium more efficiently. In this regime, the anisotropic stress responsible for GW damping is erased, leading to a pronounced enhancement of power at high multipoles relative to the free-streaming case. This behavior highlights the fluid-like character of DR at late times. In contrast, if $z_{\rm int}<z_{\rm CMB}\sim 10^{3}$, the self-interactions become efficient only after recombination and, therefore, do not affect the primary CMB anisotropies.
@@ IV Impact on the tensor-to-scalar ratio
and the white-noise weight per steradian is defined as
$w^{-1}\;=\;\frac{4\pi\,s^{2}\,f_{\rm sky}}{T_{\rm obs}}.$ (19)
Here $\theta_{\rm FWHM}$ is the full-width at half-maximum of the Gaussian beam, $s$ is the detector sensitivity (we adopt $s=1\ \mu{\rm K}\sqrt{\rm s}$), $f_{\rm sky}$ is the observed sky fraction (we adopt $f_{\rm sky}=0.1$), and $T_{\rm obs}$ is the total observation time (to be used consistently in seconds when evaluating $w^{-1}$). The bin width $\Delta\ell$ reflects our chosen multipole binning scheme (see below). The covariance is evaluated per bin and is diagonal in our setup, as indicated by the Kronecker delta $\delta_{\ell\ell^{\prime}}$.
We adopt a binning scheme that ensures an equal number of modes per bin. Concretely, we set $N_{\rm bin}=100$ bins spanning $\ell_{\rm min}=10$ to $\ell_{\rm max}=1000$, such that each bin contains the same number of harmonic modes. Within this observational model we consider two instrumental configurations to explore the dependence of our conclusions on the representative experimental sensitivity of the next-generation CMB experiments: (i) Noise 1: $\theta_{\rm FWHM}\simeq 1^{\prime}$ and $T_{\rm obs}=10$ yr; and (ii) Noise 2: $\theta_{\rm FWHM}\simeq 2^{\prime}$ and $T_{\rm obs}=5$ yr. The adopted values for $s$, $f_{\rm sky}$ and the two $\theta_{\rm FWHM}$ / $T_{\rm obs}$ combinations fully determine the noise term $w^{-1}B_{\ell}^{-2}$ entering the covariance.
@@
as a function of the fiducial value $r_{\rm fid}$. We restrict to $r_{\rm fid}<0.03$, consistent with current observational bounds [36]. The results are obtained from $500$ independent noise realizations, using a representative choice of DR parameters with $\Delta N_{\rm eff}=0.5$ and $z_{\rm int}=10^{7}$ (corresponding to $\lambda\simeq 10^{-10}$). The STD–STD and INT–INT cases are consistency checks that validate the numerical pipeline, in both the fiducial input and fitting models match and the recovered distribution of $\Delta r$ do not show any appreciable bias within statistical expectations. In contrast, the INT–STD case, for which the input fiducial cosmology contains interacting DR but the fitting model assumes standard free streaming, indicates a clear systematic overestimation of $r$, i.e. the inferred tensor amplitude is consistently biased high.
@@
Figure 4: Bias in the reconstructed tensor-to-scalar ratio $\Delta r=r_{\rm best-fit}-r_{\rm fid}$, as a function of the fiducial value of $r$, for different values of $\Delta N_{\rm eff}$, $z_{\rm int}\simeq 10^{7}$ ($\lambda\simeq 10^{-10}$), and two experimental noise configurations: Noise 1 (dark blue) and Noise 2 (yellow) representative of next-generation CMB experiments. Solid, dashed, and dotted lines correspond to $\Delta N_{\rm eff}=0.1$, $0.3$, and $0.5$, respectively. The shaded regions indicate the standard deviation related to the posterior distribution in Fig. 3 for Noise 1 (dark blue), Noise 2 (yellow), and a cosmic variance–limited scenario (dark gray). The results show that the bias can be comparable to or even exceed the statistical uncertainty, underscoring the need to accurately model DR interactions in future precision measurements of primordial B-modes. In this analysis, we isolate the impact of self-interacting DR on the estimation of $r$ by considering only this specific systematic effect, excluding others.
@@
Figure 5: Bias in the recovered tensor-to-scalar ratio, $\Delta r$, as a function of the fiducial $r$. Different line styles indicate DR self-interaction strengths, parametrized by the recoupling redshift $z_{\rm int}$ (or equivalently, the self-coupling $\lambda$). Colors correspond to DR abundance: $\Delta N_{\rm eff}=0.5$ (dark blue), $\Delta N_{\rm eff}=0.3$ (yellow), and $\Delta N_{\rm eff}=0.1$ (green). Three representative interaction regimes are shown: early ($z_{\rm int}\simeq 10^{7}$), intermediate ($z_{\rm int}\simeq 10^{4.6}$), and late ($z_{\rm int}\simeq 10^{0.9}$) recoupling. Shaded regions represent the $1\sigma$ statistical uncertainty on $r$ for the experimental configuration Noise 2 (light gray) and the cosmic variance–limited case (dark gray). As expected from the behavior shown in Fig. 2, the bias saturates for early interactions ($z_{\rm int}\gg z_{\rm CMB}\sim 10^{3}$, i.e., large $\lambda$) and it is comparable to the $1\sigma$ level, while it becomes negligible for late interactions ($z_{\rm int}\lesssim z_{\rm CMB}$, i.e. small $\lambda$), where DR interactions become efficient too late to affect the primordial B-mode signal.
@@ V Discussion and conclusions
Our results show that the systematic bias in the recovered tensor amplitude can be non-negligible. For parameter choices near current upper bounds (e.g., $\Delta N_{\rm eff}\sim 0.1$$0.5$ and early recoupling), the inferred $r$ may be overestimated by an amount comparable to, or even exceeding, the statistical sensitivities forecasted for forthcoming CMB experiments. The projected sensitivities at the level $\sigma(r)\sim 10^{-3}$ for experiments such as Simons Observatory and LiteBIRD, and the still more ambitious targets ($\sigma(r)\lesssim O(1)\times 10^{-4}$) envisaged for concepts like CMB-S4/PICO [4, 37, 21, 22, 23, 38, 39, 37, 24], highlight the observational relevance of this bias in the near future. Figures 4 and 5 illustrate this result by showing that, for moderate tensor amplitudes $r\gtrsim 5\times 10^{-3}$, the bias $\Delta r$ can reach the level of representative noise across different values of $\Delta N_{\rm eff}$, $\lambda$ (or $z_{\rm int}$), and experimental configurations.
@@
From the standpoint of data analysis, our results motivate the inclusion of parametrized models of DR self-interactions in likelihood pipelines used to infer $r$; to jointly constrain $\Delta N_{\rm eff}$ and interaction parameters rather than assuming a free-streaming DR model; and quantify the sensitivity of $r$ constraints to unmodeled interactions as part of systematic-error budgets for next-generation surveys.
Future work should extend our study in several directions. It will be important to explore a wider class of interaction operators and particle models (fermionic DR, scalar and vector mediators), study degeneracies with foreground modeling and delensing residuals in realistic pipeline-level forecasts, and assess complementarity with non-CMB probes, such as Big Bang nucleosynthesis, large-scale structure, and laboratory/astrophysical searches, that can independently constrain $\Delta N_{\rm eff}$ and interaction strengths. Addressing these points will refine the criteria for when DR interactions must be modeled to avoid biased inferences about primordial GW.

### v2

@@ I Introduction
It is well known that the coupling between ultra-relativistic (UR) species present in the primordial plasma and the GW modifies the primordial $B$-mode spectrum of the CMB polarization [56, 67, 66], thereby affecting the observable tensor-to-scalar ratio today. UR species include photons and those non-photonic relativistic components commonly grouped as neutrinos and possible beyond-Standard-Model relics [31, 33, 36], referred to as dark radiation (DR). After decoupling from the plasma, UR species are typically treated as free-streaming: in that limit they develop a nonvanishing anisotropic stress which back-reacts on the GW evolution producing a damping of the amplitude of the GW spectrum. Conversely, when interactions among UR species are efficient they isotropize the distribution function, suppress anisotropic stress and thus remove the GW damping. The competition between the species’ interaction rate and the Hubble expansion rate therefore determines whether the species behave as a free-streaming gas or as a fluid-like component, with direct consequences for the amplitude and shape of the primordial GW spectrum [50, 51, 15, 48]. The current GW spectrum including these effects was explicitly computed in [51, 48], while the effect on the primordial $B$-mode spectrum is studied in this paper.
The key question motivating this work is whether self-interactions among UR species, and in particular among a cosmologically relevant DR component, can alter the observed $B$-mode signal sufficiently to bias the inference of $r$. While current CMB and large-scale-structure measurements place strong limits on many classes of neutrino interactions at the epochs probed by the primary CMB anisotropies [62, 27, 58, 24, 8, 57, 39, 11, 23, 32, 55, 45], models of new self-interacting DR remain phenomenologically viable. In these scenarios the DR abundance, parametrized by $\Delta N_{\rm eff}$, is tightly constrained by data [6, 25, 22, 9, 59], whereas the possible types of interactions and their strengths can still cover a broad parameter space without being excluded [19, 64, 14, 20, 60, 34, 18, 21, 12, 26].
@@
The current experimental landscape of CMB polarization includes imaging telescopes such as BICEP/Keck [41, 4], CLASS [29], ACT [47], and SPT-3G [30], which are already delivering high-quality with the most stringent constraint on the tensor-to-scalar ratio to date $r<0.032$ at 95% CL by BICEP/Keck and Planck PR4 [65, 4]. QUBIC [37, 52], as the first bolometric interferometer telescope, provides a complementary approach to these measurements. Upcoming experiments, such as The Simons Observatory (SO) and the LiteBIRD satellite, are designed to reach sensitivities to primordial $B$-modes at the level of $r\sim 10^{-3}$. SO will deploy three Small Aperture Telescopes (SATs) and one Large Aperture Telescope (LAT) in the Atacama Desert, with a total of nearly 60,000 TES bolometers spanning six frequencies from 27 to 280 GHz [5, 2]. The SATs will survey approximately 10% of the sky with a projected white noise level of $2\,\mu{\rm K}$-arcmin in combined 93/145 GHz bands, enabling a constraint on the tensor-to-scalar ratio with an expected sensitivity of $\sigma(r)\simeq 0.003$. The LAT will cover about 40% of the sky with arcminute resolution and $6\,\mu{\rm K}$-arcmin noise in the same bands. Complementing these ground-based measurements, LiteBIRD (scheduled for launch to L2 around 2032) will perform a full-sky survey over three years with 15 frequency bands between 34 and 448 GHz. Its typical angular resolution is $\sim 0.5^{\circ}$ at 100 GHz, with a target sensitivity of $2.2\,\mu{\rm K}$-arcmin in polarization, designed to achieve $\delta r\simeq 0.001$ including systematic uncertainties [10, 35]. Together, the deep and high-resolution observations of SO and the all-sky coverage and systematic control of LiteBIRD will provide a decisive test of inflationary models. Beyond SO and LiteBIRD, the Probe of Inflation and Cosmic Origins (PICO) is a NASA probe-class mission concept designed to map the full sky with an unprecedented sensitivity of $0.61\,\mu{\rm K}$-arcmin across 21–799 GHz, enabling constraints on inflationary physics, neutrino masses, structure formation, and Galactic foregrounds [69, 38, 13].
@@ II.1 First-order perturbations
The first-order Einstein–Boltzmann system couples the metric to the energy–momentum tensor through Einstein equations, and the one-particle distribution function of the different species to the collision integral through the Boltzmann equations. Since our main interest is the impact on primordial CMB $B$-modes, it is sufficient to focus only on the evolution of the tensor sector within the scalar–vector–tensor (SVT) decomposition of cosmological perturbations, assuming a homogeneous and isotropic Friedmann–Lemaître–Robertson–Walker (FLRW) background [61, 42, 43, 68].
@@
For the Boltzmann equations, we include the dynamics of the one-particle distribution function of DR can be written as
@@
Restricting to the tensor sector ($m=2$), the Boltzmann hierarchy becomes {align} Θ’^(2)_2 = -57 k Θ^(2)_3 - α_2 a ⟨Γ⟩Θ^(2)_2 - ˙h,
Θ’^(2)_ℓ= k [ s2ℓ+1Θ^(2)_ℓ-1 - sℓ+12ℓ+3Θ^(2)_ℓ+1 ] - α_ℓa ⟨Γ⟩Θ^(2)_ℓ, with $s_{\ell}=\sqrt{(\ell^{2}-4)/\ell^{2}}$. The hierarchy is closed using the standard truncation described in [49], as implemented in CLASS. The relaxation-time coefficients $\alpha_{\ell}$ were computed in [19]; for simplicity, we adopt $\alpha_{\ell}=\alpha_{2}\simeq 0.188$.
@@
Finally, we implement Eqs. (11), (13), and (14) into the Einstein–Boltzmann solver CLASS [17, 46] [^1] In practice, CLASS formulates the equations using a different variable, $F^{(m)}_{\ell}$, to describe the perturbations of the one-particle distribution function. The relation between $F^{(m)}_{\ell}$ and $\Theta^{(m)}_{\ell}$ is detailed in [63]. to compute the evolution of cosmological perturbations.
@@ III Impact on $C^{BB}_{\ell}$
The main effect of these interactions is an effective enhancement of the $B$-mode power spectrum amplitude across all multipoles in comparison to the standard free-streaming DR case, with the exception of $\ell\sim 100$. This enhancement grows with increasing coupling constant $\lambda$ (or, equivalently, higher $z_{\rm int}$). Earlier transitions (larger $\lambda$ or $z_{\rm int}$) correspond to stronger self-interactions that drive the one-particle distribution function of DR toward equilibrium more efficiently. In this regime, the anisotropic stress responsible for GW damping is erased, leading to a pronounced enhancement of power at high multipoles relative to the free-streaming case. This behavior highlights the fluid-like character of DR at late times. In contrast, if $z_{\rm int}<z_{\rm CMB}\sim 10^{3}$, the self-interactions become efficient only after recombination and, therefore, do not affect the primary CMB anisotropies.
@@ IV Impact on the tensor-to-scalar ratio
[…]
@@
as a function of the fiducial value $r_{\rm fid}$. We restrict to $r_{\rm fid}<0.03$, consistent with current observational bounds [4, 65] (see also [22]). The results are obtained from $500$ independent noise realizations, using a representative choice of DR parameters with $\Delta N_{\rm eff}=0.5$ and $z_{\rm int}=10^{7}$ (corresponding to $\lambda\simeq 10^{-10}$). The STD–STD and INT–INT cases are consistency checks that validate the numerical pipeline, in both the fiducial input and fitting models match and the recovered distribution of $\Delta r$ do not show any appreciable bias within statistical expectations. In contrast, the INT–STD case, for which the input fiducial cosmology contains interacting DR but the fitting model assumes standard free streaming, indicates a clear systematic overestimation of $r$, i.e. the inferred tensor amplitude is consistently biased high.
@@
Figure 4: Bias in the reconstructed tensor-to-scalar ratio $\Delta r=r_{\rm best-fit}-r_{\rm fid}$, as a function of the fiducial value of $r$, for different values of $\Delta N_{\rm eff}$, $z_{\rm int}\simeq 10^{7}$ ($\lambda\simeq 10^{-10}$), and two experimental noise configurations: Noise 1 (dark blue) and Noise 2 (yellow) representative of next-generation CMB experiments. Solid, dashed, and dotted lines correspond to $\Delta N_{\rm eff}=0.1$, $0.3$, and $0.5$, respectively. The shaded regions indicate the standard deviation related to the posterior distribution in Fig. 3 for Noise 1 (dark blue), Noise 2 (yellow), and a cosmic variance–limited scenario (only CV in dark gray). The results show that the bias can be comparable to or even exceed the statistical uncertainty, underscoring the need to accurately model DR interactions in future precision measurements of primordial B-modes. In this analysis, we isolate the impact of self-interacting DR on the estimation of $r$ by considering only this specific systematic effect, excluding others.
@@
Figure 5: Bias in the recovered tensor-to-scalar ratio, $\Delta r$, as a function of the fiducial $r$. Different line styles indicate DR self-interaction strengths, parametrized by the recoupling redshift $z_{\rm int}$ (or equivalently, the self-coupling $\lambda$). Colors correspond to DR abundance: $\Delta N_{\rm eff}=0.5$ (dark blue), $\Delta N_{\rm eff}=0.3$ (yellow), and $\Delta N_{\rm eff}=0.1$ (green). Three representative interaction regimes are shown: early ($z_{\rm int}\simeq 10^{7}$), intermediate ($z_{\rm int}\simeq 10^{4.6}$), and late ($z_{\rm int}\simeq 10^{0.9}$) recoupling. Shaded regions represent the $1\sigma$ statistical uncertainty on $r$ for the experimental configuration Noise 1 (light gray) and the cosmic variance–limited case (only CV in dark gray). As expected from the behavior shown in Fig. 2, the bias saturates for early interactions ($z_{\rm int}\gg z_{\rm CMB}\sim 10^{3}$, i.e., large $\lambda$) and it is comparable to the $1\sigma$ level, while it becomes negligible for late interactions ($z_{\rm int}\lesssim z_{\rm CMB}$, i.e. small $\lambda$), where DR interactions become efficient too late to affect the primordial B-mode signal.
@@ V Discussion and conclusions
Our results show that the systematic bias in the recovered tensor amplitude can be non-negligible. For parameter choices near current upper bounds (e.g., $\Delta N_{\rm eff}\sim 0.1$$0.5$ and early recoupling), the inferred $r$ may be overestimated by an amount comparable to, or even exceeding, the statistical sensitivities forecasted for forthcoming CMB experiments. The projected sensitivities at the level $\sigma(r)\sim 10^{-3}$ for experiments such as The Simons Observatory and LiteBIRD, and the still more ambitious targets ($\sigma(r)\lesssim\mathcal{O}(1)\times 10^{-4}$) envisaged for concepts like CMB-S4/PICO [3, 1, 5, 2, 10, 69, 16, 28, 1, 38, 13, 35], highlight the observational relevance of this bias in the near future. Figures 4 and 5 illustrate this result by showing that, for moderate tensor amplitudes $r\gtrsim 5\times 10^{-3}$, the bias $\Delta r$ can reach the level of representative noise across different values of $\Delta N_{\rm eff}$, $\lambda$ (or $z_{\rm int}$), and experimental configurations.
From the standpoint of data analysis, our results motivate the inclusion of parametrized models of DR self-interactions in likelihood pipelines used to infer $r$; to jointly constrain $\Delta N_{\rm eff}$ and interaction parameters rather than assuming a free-streaming DR model; and quantify the sensitivity of $r$ constraints to unmodeled interactions as part of systematic-error budgets for next-generation surveys. In particular, the bias estimated in this work, $\Delta r\gtrsim 10^{-4}$, exceeds the LiteBIRD budget allocation per systematic effect ($\Delta r_{\rm sysLB}\simeq 6.5\times 10^{-6}$[35].
@@
As future work, it will be important to explore a wider class of interaction operators and particle models (fermionic DR, scalar and vector mediators), study degeneracies with foreground modeling and delensing residuals in realistic pipeline-level forecasts, and assess complementarity with non-CMB probes, such as Big Bang nucleosynthesis, large-scale structure, and laboratory/astrophysical searches, that can independently constrain $\Delta N_{\rm eff}$ and interaction strengths. Addressing these points will refine the criteria for when DR interactions must be modeled to avoid biased inferences about primordial GW.